OpenStudy (anonymous):

As the visiting year 10 hockey team were first to arrive at the changing rooms, they had their choice of lockers. Every locker is given a number in order and they are arranged in three rows, numbered 1, 2, 3, ... on the top row. The sequence continuing on the next row. Four friends quickly claimed the corner lockers, Rather amazingly the four friends found that the numbers on their lockers added up to the sum of their ages. How many lockers are there? Note: In England, a year 10 student will be 14 or 15 years old.

OpenStudy (anonymous):

Well the locker numbers would add up to \[1+x+(2x+1)+3x\] assuming that x=the number of lockers in the top row

OpenStudy (anonymous):

So that's \[6x+2\] and this could be equal to 56, 57, 58 or 59

OpenStudy (anonymous):

There is only one number that comes out even, 56, where \[6x+2=56\] \[6x=54\] \[x=9\]

OpenStudy (anonymous):

have u looked at the xmas tree one?

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

When Christmas trees are planted they should stand at least 2 metres away from one another whilst growing. What is the maximum number of trees that can be planted in one square kilometre?

OpenStudy (anonymous):

Rows of height \[\sqrt{3}\] So 577 rows, alternating 500 and 499 trees So that's 144500+143712=288212 trees, this is my final answer :)

OpenStudy (anonymous):

I think it is right :-)

OpenStudy (anonymous):

yay :)

OpenStudy (anonymous):

288673

OpenStudy (anonymous):

aww damn

OpenStudy (anonymous):

That's a lot of trees I missed

OpenStudy (anonymous):

1000000/2sqrt3